OpenStudy (anonymous):

Line passes through the points (1, 4) and (4, 9). Line m passes through the point (3, 2) and is parallel to line . What is the equation for line m?

OpenStudy (anonymous):

The first line has slope \[\text{slope}=\frac{9-4}{4-1}=\frac{5}{3}.\]The second line will as well have the same slope m = 5/3 since the two lines are parallel. [Now use the point-slope formula, using (3, 2), to find the line m.]